# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ 5.2 Polarization Comparison with Dwingeloo
The polarized intensity in DRAGONS is consistently lower than in Dwingeloo, and the discrepancy is more pronounced at the lower frequencies. This is likely caused by more beam depolarization present in DRAGONS due to its beam having approximately twice the area of the Dwingeloo beam. Since the amount of beam depolarization does not only depend on the size of the beam, but also on the angular scale of the fluctuations in Stokes $Q$ and $U$, it is not possible to determine the expected difference in beam depolarization between the two surveys without comparing to a higher resolution dataset. Once the full-sky Galactic foreground polarization data from CHIME are calibrated, the CHIME Stokes $Q$ and $U$ maps, at $\sim 20\arcmin$ resolution, can provide the necessary small-scale information for this calculation.
Figure 28: Comparisons between DRAGONS and Dwingeloo polarization data for the four frequencies common to both surveys (rows ad). The first two columns show Stokes $Q$ and $U$ comparisons, the third column shows PI, and the fourth column shows polarization angles. Slopes, intercepts, and correlation coefficients are shown for all four quantities. For Stokes $Q$ and $U$ we exclude data points with PI below three times the noise level from the linear fit calculations to avoid biasing the fits with low signal values. The generally lower polarized intensities for DRAGONS compared to Dwingeloo can be attributed to the lower angular resolution of DRAGONS resulting in more beam depolarization.
5.3 Error Analysis
@@
Figure 29: Polarized intensity (ac) and systematic error in Stokes $Q$ and $U$ (df) for three sample frequency channels. The spatially-dependent, systematic error, $\delta_{QU_{\text{sys}}}$, is calculated as the average of the errors in $Q$ and $U$, which are in turn calculated as $\delta Q=|Q_{E}-Q_{W}|$ and $\delta U=|U_{E}-U_{W}|$. These error estimates reflect both spatial and spectral variation in the reliability of the polarization maps based on the differences between the independent East and West maps that are ultimately combined. We combine these systematic errors in quadrature with the measured random spatial variability to provide uncertainty maps for Faraday synthesis. The grey regions correspond to the four brightest radio sources, which were masked out in the map-making process.
Particularly at frequencies below $\sim 500$ MHz, residual errors following the ground and ionosphere correction steps (see Section 4.6) lead to the East and West maps differing over some regions of the sky. At each frequency, we use the differences between the East and West maps as a proxy for the resulting position-dependent, systematic error in Stokes $Q$ and $U$, calculated as
@@
where $Q_{\text{E}}$ ($U_{\text{E}}$) is the Stokes $Q$ ($U$) East map and $Q_{\text{W}}$ ($U_{\text{W}}$) is the Stokes $Q$ ($U$) West map. Examples of the resulting error maps are shown in Figure 29, compared to the polarized intensity maps. Instrumental polarization leakage leads to enhanced errors along the brightest part of the Galactic disk (RA=18 to 22 hours). This is present across the full band, generally making the area within approximately two beamwidths of the Galactic plane toward the inner Galaxy unreliable. Enhanced errors resulting from the residual systematics (inadequately corrected ground and ionosphere effects) occur mostly at low frequencies. In particular, the errors are enhanced in patches of the sky that were observed under significantly different ionospheric conditions between East and West maps.
To quantify the random spatial variability (noise) in the maps, we calculated the standard deviation in Stokes $Q$ and $U$ in a region of low polarized intensity in the longitude range $180\arcdeg\leq\ell\leq 240\arcdeg$. These values range from 0.5 to 0.05 K across 350 to 1030 MHz. We combine the two types of errors described here in quadrature, and use the resulting error cubes (with both spatial and spectral dependence) as inputs to Faraday synthesis (Section 6) for determining position-dependent channel weighting.
@@
5.4 Polarization Leakage
For circular feeds, imperfections in the correlated outputs, RL* and LR*, can lead to on-axis leakage of Stokes $I$ into $Q$ and $U$. This leakage is almost completely removed by the daily median subtraction (Step 2 in Section 4.3). We examine the final Stokes $I$, $Q$, and $U$ maps to quantify the amount of residual leakage. The Orion A H II region ($\ell=209\arcdeg$, $b=-19\arcdeg$), at $\sim 1\arcdeg$ in angular size, is unresolved at most DRAGONS frequencies and produces the expected cloverleaf pattern characterizing the DRAO-15 beam pattern in Stokes $Q$ and $U$ (Figure 23 and 24). At its center, we measure less than 1% fractional polarization across most of band, increasing to 2% at the higher frequencies where the source may be partially resolved. This allows us to identify an upper limit of 1–2% on-axis polarization leakage across the DRAGONS band.
The Galactic midplane stands out prominently as a band of instrumental polarization in the Stokes $Q$, $U$, PI, and polarization angle maps (Figures 2326) at Galactic longitudes $\ell<90\arcdeg$. The midplane should be depolarized in the DRAGONS data due to polarization structures in the Galactic disk on scales much smaller than the beam. Although the fractional polarization in the midplane does not differ significantly from the surroundings in our maps, there is a significant increase in PI within two beamwidths on either side of the Galactic midplane for $\ell<90\arcdeg$. Therefore, $\pm 7\arcdeg$ in latitude from the midplane at 350 MHz, and $\pm 2\arcdeg$ at 1030 MHz should be excluded from analysis. For Faraday synthesis we mask $\pm 7\arcdeg$ since the beam is convolved to the common resolution of the beam at 350 MHz and all frequencies are included in calculating the spectra.
@@ 5.4 Polarization Leakage
Noise, $Q$ and $U$ images (single channel) 500 to 60 mK
Available data formats Galactic coordinates, FITS and HEALPix
@@
Largest measurable Faraday depth structure 38 rad m-2
@@
6 Faraday Depth Cubes
@@
The primary data product from the DRAGONS survey is a diffuse-emission Faraday depth cube, which can be studied spectrally along individual lines of sight, or in 2D maps, either as Faraday depth slices or as moments of the Faraday spectra (Dickey et al., 2019). The strength of the DRAGONS survey is that the frequency range allows for high resolution in Faraday depth with simultaneous sensitivity to broad Faraday depth structures. The Faraday depth resolution, $\delta\phi$, is inversely proportional to the $\lambda^{2}$ coverage (Brentjens & de Bruyn, 2005), and is 6 rad m-2 for DRAGONS. The widest detectable (not depolarized) feature in Faraday depth space, $\phi_{\mathrm{max-scale}}$, is inversely proportional to the smallest $\lambda^{2}$ in the survey. For DRAGONS this is 38 rad m-2. Diffuse emission Faraday synthesis has the potential to detect the widest range of scales of Faraday depth structures when $\phi_{\mathrm{max-scale}}>\delta\phi$. DRAGONS is the only GMIMS component survey to date to achieve this criterion by a significant margin, resulting in DRAGONS having unprecedented sensitivity to diffuse emission Faraday complexity. The published data parameters are summarized in Table 3 and the available data products are listed in Table 4.
6.1 Faraday Synthesis
@@
For applying Faraday synthesis, we convolved the 1 MHz binned frequency cubes of Stokes $Q$, $U$, and $I$, with persistent RFI channels flagged, to a common angular resolution of 3.6°(the survey resolution at 350 MHz). In addition to the flagged RFI channels, we also conducted a visual inspection to mask out either entire maps or portions of maps in channels heavily contaminated by other artifacts. We then fed the remaining set of 393 full maps and 11 partial maps into the CIRADA RMTools package (Purcell et al., 2020) [^12] https://github.com/CIRADA-Tools/RM-Tools for Faraday synthesis and RM CLEAN calculations.
@@
For Faraday synthesis, we used inverse variance weighting of the channels to account for the higher noise and systematic errors at lower frequencies. This was based on the error maps described in Section 5.3. We incorporated Stokes $I$ data (corrected for the zero-offset as described in Section 5.1) as a fitted model along each LOS to account for the spectral index variation across the 680 MHz-wide band. The effect of dividing by the Stokes $I$ model (and thus using fractional polarization as the input to Faraday synthesis) was to reduce the uncertainties at the lower frequencies and thus partially offset the down-weighting of the low-frequency channels introduced by the inverse variance weighting. Examples of the resulting rotation measure spread function (RMSF) are shown in Figure 30.
@@
Figure 31: Sample Faraday depth spectra, corresponding to the frequency spectra shown in Figure 21. The dirty (cleaned) spectra are shown in grey (black), the CLEAN threshold is shown by the green dot-dashed line, the minimum PI threshold for the moment calculations is the pink dashed line, and the CLEAN components are in blue.
6.2 Faraday Moment Maps
Following previous GMIMS work, such as Dickey et al. (2019) and Raycheva et al. (2025), we produce Faraday depth moment maps in order to collapse the three-dimensional data cubes into 2D maps for ease of visualization and for studying the large-scale plane-of-sky structures in the Faraday rotation. The zeroth moment (M0; Figure 32) is the total polarized intensity integrated over a selected Faraday depth range, the first moment (M1; Figure 33) is the polarized-intensity-weighted average of the Faraday depths, and the square root of the second moment (m2; Figure 34) [^13] We use ‘m2’ (units of rad m-2) to refer to the square root of the second moment, M2. represents the width of the distribution of Faraday depths. The equations defining these metrics are given in Dickey et al. (2019). For the maps presented in Figures 3234, we calculated the moments on the cleaned Faraday depth cube out to $\pm 100$ rad m-2, using a threshold of $6\sigma$ (derived from the dirty Faraday depth cubes) as the minimum polarized intensity to include in the calculations. This is the first diffuse emission polarization survey that has the sensitivity to compute Faraday moments almost everywhere, with only a small blank patch (near $\ell=30\arcdeg,b=-15\arcdeg$) where the PI is too low to give a clear detection in the Faraday spectrum.
@@ 7.1 Faraday Complexity
We study the effect further by counting the number of distinct restored Faraday depth peaks above $8\sigma$ along each LOS and mapping the number of peak, as shown in Figure 35. We note that the Fan Region, ($\ell\approx 130\arcdeg$, $b\approx 5\arcdeg$) is one of the few large-scale Faraday simple regions (having a single Faraday depth peak) in the northern Galactic hemisphere, while in the southern hemisphere there is a larger area of Faraday simple lines of sight. The North Polar Spur (NPS; Sun et al., 2015a), near $\ell\approx 20\arcdeg$, $b>25\arcdeg$, and the North Celestial Pole loop (NCPL; Meyerdierks et al., 1991; Marchal & Martin, 2023), near $\ell\approx 30\arcdeg$, $120\arcdeg<b<150\arcdeg$, are two of the regions exhibiting the most Faraday complexity in the DRAGONS data, likely indicating multiple emission and rotation regions along the lines of sight. The second and third panels of Figure 31 show sample Faraday depth spectra from the NPS and NCPL respectively. The spread of the features in Faraday depth space being comparable to the widest detectable Faraday depth scale of $\phi_{\mathrm{max-scale}}=38$ rad m-2 is also consistent with partially depolarized extended features in Faraday depth resulting from mixed emission and rotation. Modeling these lines of sight using methods like ‘$QU$-fitting’ (Farnsworth et al., 2011; O’Sullivan et al., 2012; Ideguchi et al., 2014; Sun et al., 2015b) will be enlightening in revealing the ISM geometry that produces the complex structures.
@@
7.2 Comparison with GMIMS-LBS
DRAGONS and its southern-hemisphere counterpart, GMIMS-LBS, overlap in frequency between 350 MHz and 480 MHz, and in declination coverage from $\delta=-20\arcdeg$ to $\delta=20\arcdeg$, corresponding to 34% of the sky in solid angle. This makes the surveys ideally suited for comparison and GMIMS data-product validation. Given the 64 m diameter of the Murriyang telescope (used for all southern hemisphere GMIMS surveys), DRAGONS has approximately four times the beamwidth of GMIMS-LBS. Thus, prior to comparing, we convolve the GMIMS-LBS maps to the DRAGONS beam. We show a point-by-point comparison on a HEALPix grid with nside=64 for Stokes $Q$ and $U$ at 408 MHz in Figure 36, masking out a region $\pm 10\arcdeg$ around the Galactic plane, where both data sets suffer from instrumental polarization leakage. Both Stokes $Q$ and $U$ are strongly correlated between the two surveys with Pearson correlation coefficients of 0.63, and have ratios close to unity, indicating successful calibration of both surveys.

### v2

@@ V.2 Polarization Comparison with Dwingeloo
The PI in DRAGONS is consistently lower than in Dwingeloo, and the discrepancy is more pronounced at the lower frequencies. This is likely caused by more beam depolarization present in DRAGONS due to its beam having approximately twice the area of the Dwingeloo beam. Since the amount of beam depolarization depends not only on the size of the beam, but also on the angular scale of the fluctuations in Stokes $Q$ and $U$, it is not possible to determine the expected difference in beam depolarization between the two surveys without comparing to a higher resolution dataset. Once the full-sky Galactic foreground polarization data from CHIME are calibrated, the CHIME Stokes $Q$ and $U$ maps, at $\sim 20\arcmin$ resolution, can provide the necessary small-scale information for this calculation.
Figure 28: Comparisons between DRAGONS and Dwingeloo polarization data for the four frequencies common to both surveys (rows ad). The first two columns show Stokes $Q$ and $U$ comparisons, the third column shows PI, and the fourth column shows polarization angles. Slopes, intercepts, and correlation coefficients are shown for all four quantities. For Stokes $Q$ and $U$ we exclude data points with PI below 3 times the noise level from the linear fit calculations to avoid biasing the fits with low signal values. The generally lower polarized intensities for DRAGONS compared to Dwingeloo can be attributed to the lower angular resolution of DRAGONS resulting in more beam depolarization.
V.3 Error Analysis
@@
Figure 29: Polarized intensity (ac) and systematic error in Stokes $Q$ and $U$ (df) for three sample frequency channels. The spatially-dependent, systematic error, $\delta_{QU_{\text{sys}}}$, is calculated as the average of the errors in $Q$ and $U$, which are in turn calculated as $\delta Q=|Q_{E}-Q_{W}|$ and $\delta U=|U_{E}-U_{W}|$. These error estimates reflect both spatial and spectral variation in the reliability of the polarization maps based on the differences between the independent east and west maps that are ultimately combined. We combine these systematic errors in quadrature with the measured random spatial variability to provide uncertainty maps for Faraday synthesis. The gray regions correspond to the four brightest radio sources, which were masked out in the map-making process.
Particularly at frequencies below $\sim 500$ MHz, residual errors following the ground and ionosphere correction steps (see Section IV.6) lead to the east and west maps differing over some regions of the sky. At each frequency, we use the differences between the east and west maps as a proxy for the resulting position-dependent, systematic error in Stokes $Q$ and $U$, calculated as
@@
where $Q_{\text{E}}$ ($U_{\text{E}}$) is the Stokes $Q$ ($U$) east map and $Q_{\text{W}}$ ($U_{\text{W}}$) is the Stokes $Q$ ($U$) west map. Examples of the resulting error maps are shown in Figure 29, compared to the PI maps. Instrumental polarization leakage leads to enhanced errors along the brightest part of the Galactic disk (RA=18 to 22 hours). This is present across the full band, generally making the area within approximately two beamwidths of the Galactic plane toward the inner Galaxy unreliable. Enhanced errors resulting from the residual systematics (inadequately corrected ground and ionosphere effects) occur mostly at low frequencies. In particular, the errors are enhanced in patches of the sky that were observed under significantly different ionospheric conditions between east and west maps.
To quantify the random spatial variability (noise) in the maps, we calculated the standard deviation in Stokes $Q$ and $U$ in a region of low polarized intensity in the longitude range $180\arcdeg\leq\ell\leq 240\arcdeg$. These values range from 0.5 to 0.05 K across 350 to 1030 MHz. We combine the two types of errors described here in quadrature, and use the resulting error cubes (with both spatial and spectral dependence) as inputs to Faraday synthesis (Section VI) for determining position-dependent channel weighting.
@@
V.4 Polarization Leakage
For circular feeds, imperfections in the correlated outputs, RL* and LR*, can lead to on-axis leakage of Stokes $I$ into $Q$ and $U$. This leakage is almost completely removed by the daily median-subtraction (Step 2 in Section IV.3). We examine the final Stokes $I$, $Q$, and $U$ maps to quantify the amount of residual leakage. The Orion A H II region ($\ell=209\arcdeg$, $b=-19\arcdeg$), at $\sim 1\arcdeg$ in angular size, is unresolved at most DRAGONS frequencies and produces the expected cloverleaf pattern characterizing the DRAO-15 beam pattern in Stokes $Q$ and $U$ (Figure 23 and 24). At its center, we measure less than 1% fractional polarization across most of band, increasing to 2% at the higher frequencies where the source may be partially resolved. This allows us to identify an upper limit of 1–2% on-axis polarization leakage across the DRAGONS band.
The Galactic midplane stands out prominently as a band of instrumental polarization in the Stokes $Q$, $U$, PI, and polarization angle maps (Figures 2326) at Galactic longitudes $\ell<90\arcdeg$. The midplane should be depolarized in the DRAGONS data due to polarization structures in the Galactic disk on scales much smaller than the beam. Although the fractional polarization in the midplane does not differ significantly from the surroundings in our maps, there is a significant increase in PI within two beamwidths on either side of the Galactic midplane for $\ell<90\arcdeg$. Therefore, $\pm 7\arcdeg$ in latitude from the midplane at 350 MHz, and $\pm 2\arcdeg$ at 1030 MHz should be excluded from analysis. For interpreting Faraday synthesis results we recommend masking $\pm 7\arcdeg$ since the beam is convolved to the common resolution of the beam at 350 MHz and all frequencies are included in calculating the spectra.
@@ V.4 Polarization Leakage
Noise, $Q$ and $U$ images 500 to 60 mK
(single channel)
Available data formats Galactic coordinates,
FITS and HEALPix
@@
Largest measurable 38 rad m-2
Faraday depth structure
@@
Number of Faraday depth peaks
VI Faraday Depth Cubes
@@
The primary data product from the DRAGONS survey is a diffuse-emission Faraday depth cube, which can be studied spectrally along individual lines of sight, or in 2D maps, either as Faraday depth slices or as moments of the Faraday spectra (Dickey et al., 2019). The strength of the DRAGONS survey is that the frequency range allows for high resolution in Faraday depth with simultaneous sensitivity to broad Faraday depth structures. The Faraday depth resolution, $\delta\phi$, is inversely proportional to the $\lambda^{2}$ coverage (Brentjens and de Bruyn, 2005), and is 6 rad m-2 for DRAGONS. The widest detectable (not completely depolarized) feature in Faraday depth space, $\phi_{\mathrm{max-scale}}$, is inversely proportional to the smallest $\lambda^{2}$ in the survey. For DRAGONS this is 38 rad m-2 based on Brentjens and de Bruyn (2005) [^12] A more conservative estimate of this quantity is presented in Rudnick and Cotton (2023). Diffuse emission Faraday synthesis has the potential to detect the widest range of scales of Faraday depth structures when $\phi_{\mathrm{max-scale}}>\delta\phi$. DRAGONS is the only GMIMS component survey to date to achieve this criterion by a significant margin, resulting in DRAGONS having unprecedented sensitivity to diffuse emission Faraday complexity. The published data parameters are summarized in Table 3 and the available data products are listed in Table 4.
VI.1 Faraday Synthesis
@@
For applying Faraday synthesis, we convolved the 1 MHz binned frequency cubes of Stokes $Q$, $U$, and $I$, with persistent RFI channels flagged, to a common angular resolution of 3.6°(the survey resolution at 350 MHz). In addition to the flagged RFI channels, we also conducted a visual inspection to mask out either entire maps or portions of maps in channels heavily contaminated by other artifacts. We then fed the remaining set of 393 full maps and 11 partial maps (out of a total of 680) into the CIRADA RMTools package (Purcell et al., 2020) [^13] https://github.com/CIRADA-Tools/RM-Tools for Faraday synthesis and RM CLEAN calculations.
Figure 31: Sample Faraday depth spectra, corresponding to the frequency spectra shown in Figure 21. The dirty (cleaned) spectra are shown in gray (black), the CLEAN threshold is shown by the green dot-dashed line, the minimum PI threshold for the moment calculations is the pink dashed line, and the CLEAN components are in blue.
@@
For Faraday synthesis, we used inverse variance weighting of the channels to account for the higher noise and systematic errors at lower frequencies. This was based on the error maps described in Section V.3. We incorporated Stokes $I$ data (corrected for the zero-offset as described in Section V.1) as a fitted model along each LOS to account for the spectral index variation across the 680 MHz-wide band. The effect of dividing by the Stokes $I$ model (and thus using fractional polarization as the input to Faraday synthesis) was to reduce the uncertainties at the lower frequencies and thus partially offset the down-weighting of the low-frequency channels introduced by the inverse variance weighting. Examples of the resulting rotation measure spread function (RMSF) are shown in Figure 30. The investigation of alternative weighting schemes will be the subject of future studies.
@@
VI.2 Faraday Moment Maps
Following previous GMIMS work, such as Dickey et al. (2019) and Raycheva et al. (2025), we produce Faraday depth moment maps in order to collapse the three-dimensional data cubes into 2D maps for ease of visualization and for studying the large-scale plane-of-sky structures in the Faraday rotation. The zeroth moment (M0; Figure 32) is the total polarized intensity integrated over a selected Faraday depth range, the first moment (M1; Figure 33) is the PI-weighted average of the Faraday depths, and the square root of the second moment (m2; Figure 34) [^14] We use ‘m2’ (units of rad m-2) to refer to the square root of the second moment, M2. represents the width of the distribution of Faraday depths. The equations defining these metrics are given in Dickey et al. (2019). For the maps presented in Figures 3234, we calculated the moments on the cleaned Faraday depth cube out to $\pm 100$ rad m-2, using a threshold of $6\sigma$ (derived from the dirty Faraday depth cubes) as the minimum polarized intensity to include in the calculations. This is the first diffuse emission polarization survey that has the sensitivity to compute Faraday moments almost everywhere, with only a small blank patch (near $\ell=30\arcdeg,b=-15\arcdeg$) where the PI is too low to give a clear detection in the Faraday spectrum.
@@ VII.1 Faraday Complexity
We study the effect further by counting the number of distinct restored Faraday depth peaks above $8\sigma$ along each LOS and mapping the number of peak, as shown in Figure 35. We note that the Fan Region, ($\ell\approx 130\arcdeg$, $b\approx 5\arcdeg$) is one of the few large-scale Faraday simple regions (having a single Faraday depth peak) in the northern Galactic hemisphere, while in the southern hemisphere there is a larger area of Faraday simple lines of sight. The North Polar Spur (NPS; Sun et al., 2015a), near $\ell\approx 20\arcdeg$, $b>25\arcdeg$, and the North Celestial Pole loop (NCPL; Meyerdierks et al., 1991; Marchal and Martin, 2023), near $\ell\approx 30\arcdeg$, $120\arcdeg<b<150\arcdeg$, are two of the regions exhibiting the most Faraday complexity in the DRAGONS data, likely indicating multiple emission and rotation regions along the lines of sight. The second and third panels of Figure 31 show sample Faraday depth spectra from the NPS and NCPL respectively. The spread of the features in Faraday depth space being comparable to the widest detectable Faraday depth scale of $\phi_{\mathrm{max-scale}}=38$ rad m-2 is also consistent with partially depolarized extended features in Faraday depth resulting from mixed emission and rotation. Modeling these lines of sight using methods like ‘$QU$-fitting’ (Farnsworth et al., 2011; O’Sullivan et al., 2012; Ideguchi et al., 2014; Sun et al., 2015b) will be enlightening in revealing the ISM geometry that produces the complex structures.
@@
VII.2 Comparison with GMIMS-LBS
DRAGONS and its southern-hemisphere counterpart, GMIMS-LBS, overlap in frequency between 350 MHz and 480 MHz, and in declination coverage from $\delta=-20\arcdeg$ to $\delta=20\arcdeg$, corresponding to 34% of the sky in solid angle. This makes the surveys ideally suited for comparison and GMIMS data-product validation. Given the 64 m diameter of the Murriyang telescope (used for all southern-hemisphere GMIMS surveys), DRAGONS has approximately 4 times the beamwidth of GMIMS-LBS. Thus, prior to comparing, we convolve the GMIMS-LBS maps to the DRAGONS beam. We show a point-by-point comparison on a HEALPix grid with nside=64 for Stokes $Q$ and $U$ at 408 MHz in Figure 36, masking out a region $\pm 10\arcdeg$ around the Galactic plane, where both data sets suffer from instrumental polarization leakage. Both Stokes $Q$ and $U$ are strongly correlated between the two surveys with Pearson correlation coefficients of 0.63, and have ratios close to unity, indicating successful calibration of both surveys.
The strong agreement between GMIMS-LBS and DRAGONS allows for the maps to be combined into a single, full-sky image across the overlapping frequency range. We show examples of such a combined map in Figure 37, where we have averaged the maps in the overlapping declination range. This highlights the potential of the full GMIMS initiative in allowing for polarization and Faraday depth mapping of large-scale patterns simultaneously covering both northern and southern Galactic hemispheres. Examples of existing and forthcoming studies that take advantage of this are: (i) a follow-up to the analysis of the asymmetry between the northern and southern Galactic latitudes explored in Dickey et al. (2022) using GMIMS-HBN and STAPS (C. van Bergen et al., 2026, in preparation) and (ii) an analysis of the large-scale magnetic field reversal using STAPS and the higher frequencies (>500>500 MHz) of DRAGONS (Booth et al., 2026). Other future work will be able to benefit from the high Faraday depth resolution provided by GMIMS-LBS and the lower frequencies of DRAGONS to investigate Faraday complexity in the more nearby Galactic volume by using only the lower end of the DRAGONS band (<500<500 MHz).
VII.3 Perseus-Taurus Bubble